OpenStudy (anonymous):

what is the slope of the line that is perpendicular to the line through the points (0, -3) and (-2, -2)?

OpenStudy (anonymous):

Find the slope of that line using the slope equation: A perpendicular line's slope will be the negative reciprocal of that slope-- Example: If the slope of a line is 1/2, then the slope of a line perpendicular will be -2.

OpenStudy (texaschic101):

find the slope using the slope formula.. slope = (y2 - y1) / (x2 - x1) (0,-3) x1 = 0 and y1 = -3 (-2,-2) x2 = -2 and y2 = -2 now we sub slope = (-2 - (-3) / (-2 - 0) can you finish and find the slope ?

OpenStudy (texaschic101):

Once you do the math and find the slope, we want a perpendicular slope. To do this, you just " flip " the slope and change the sign. Like the above example : 1/2...to find negative reciprocal, " flip" the slope and change the sign...1/2 neg. rec. = -2/1 or just -2. So find the slope, then flip it and change the sign, and you will have the perpendicular slope.